### v1

# No Imputation of Missing Values In Tabular Data Classification Using Incremental Learning

Abstract
Tabular data sets with varying missing values are prepared for machine learning using an arbitrary imputation strategy. Synthetic values generated by imputation models often concern data stakeholders about computational complexity, data quality, and data-driven outcomes. This paper eliminates these concerns by proposing no imputation incremental learning (NIIL) of tabular data with varying missing value rates and types. The proposed method incrementally learns partitions of overlapping feature sets while using attention masks to exclude missing values from attention scoring. The average classification performance rank order across 15 diverse tabular data sets highlights the superiority of NIIL over 11 state-of-the-art learning methods with or without missing value imputations. Further experiments substantiate the robustness of NIIL against varying missing value types and rates compared to methods that involve the imputation of missing values. Our empirical analysis reveals that a feature partition size of half of the original feature space is, computation-wise and accuracy-wise, the best choice for the proposed incremental learning. The proposed method is one of the first deep learning solutions that can effectively learn tabular data without requiring the imputation of missing values.
Keywords missing values, incremental learning, tabular data, deep learning, imputation.
1 Introduction
Missing values at varying rates and types are ubiquitous in health science involving electronic health records (EHR) data. For two reasons, missing values remain critical to ensuring unbiased data and reliable data-driven outcomes. First, the state-of-the-art machine learning (ML) methods that promise precision medicine are not designed for data with missing values. Second, missing values are handled arbitrarily by excluding samples and features or imputing with model-generated synthetic values to address the first reason. Sample exclusion biases medical data with sicker patients when their healthy counterparts are known to have more missing values [1]. Feature exclusion due to high missing rates can miss potentially important predictors. In contrast, many health scientists remain skeptical of using medical data imputed with model-generated synthetic values, especially using a hard-to-interpret ML model. Synthetic values generated by models can inadvertently alter feature importance, data statistics, and correlations between features, dampening the predictive ability of data [2]. Multiple imputations using chained equations (MICE) is arguably the most popular method in health science to impute missing values. MICE estimates missing values using a chain of linear regression models for individual features with two major limitations. First, missing values of one feature may be independent of other features, making regression models ineffective. Second, the relationship between features is rarely linear in practice. In other words, data statistics and missing value patterns should guide the modeling and selection of an imputation algorithm. Unfortunately, there is no best imputation method for all data sets or missing value types [3, 4, 5]. All concerns surrounding imputation methods underscore the need for alternatives to imputation in learning data with missing values. This paper proposes No Imputation Incremental Learning (NIIL) of tabular data with missing values. NIIL can directly classify tabular data with missing values, eliminating biases and concerns surrounding data exclusions and synthetic value imputation.
1.1 Related work
Feature incremental learning (FIL) is one of the cornerstones of deep learning (DL), which is not viable in traditional ML. Although highly successful with homogeneous image features [6], incremental learning is challenging in the heterogeneous feature space of tabular data [7]. Moreover, the superiority of traditional ML in tabular data over deep learning is well documented in the literature [8]. Some recent work has achieved FIL on tabular data [7, 9, 10, 11]. A common approach to FIL trains multiple models separately to accommodate new features and aggregate model predictions through a weighting mechanism [10, 7]. Alternatively, mamba [9] and the transformer-based method [11] use a single model to train on additional features incrementally as they become available. However, no current deep learning or FIL methods have been proposed to directly learn tabular data with missing values without requiring imputation.
Figure 1: No Imputation Incremental Learning (NIIL) algorithm uses $P$ fixed-sized overlapping feature partitions to incrementally train a Feature-Tokenized Transformer (FTT). Attention masks: $M_{1}$ operates as the attention column mask and $M_{2}$ as the attention row mask to exclude missing feature values from attention scoring.
The state-of-the-art methods for handling missing values are broadly categorized into traditional and deep imputations. Traditional methods include MICE [12], missForest [13], and MLP in multiple imputations [5]). These methods use regression models to estimate missing values, assuming relationships between features. Deep imputation methods include Generative Adversarial Imputation Nets (GAIN) [14], diffusion-based imputation (Diffputer) [15], variational autoencoders [16], and denoising autoencoders [17]. Sun et al.  [18] report that GAIN has limited performance when missing values are at random (MAR) or not at random (MNAR), where MICE often outperforms GAIN. However, generative adversarial networks (GANs) can be unstable during optimization, particularly in terms of model convergence[19]. Diffusion-based methods use a simplified assumption about feature distributions (e.g., Gaussian distribution), which may not be optimal to yield the most accurate estimation of missing values. Chen et al. [20] report the superiority of GAIN over its diffusion counterparts, such as MissDiff [21] and CSDI_T [22]. These data-centric assumptions and limitations of the imputation methods explain why no single imputation model is ideal for all data sets. On the other hand, DL methods for missing value imputation have failed to attract domain applications due to high computational costs, lack of trust, and interpretability [23, 24]. In fact, several DL methods directly learn data with missing values by jointly imputing and classifying in a single framework [25, 26]. However, such an approach has the computational requirements of initializing and updating missing-value estimates with an imputation error. Joint learning of imputation and classification may bias the imputed values for given classification targets. Therefore, a DL framework capable of directly learning from observed values, excluding the contribution or imputation of missing values, would eliminate computational costs and biases due to synthetic data generation. In fact, decision tree-based ML models such as Random Forest and XGBoost [27] use “Missing Incorporated in Attributes” (MIA)[28] to directly learn data with missing values. However, traditional ML methods can be prone to overfitting high-dimensional feature space when incremental learning is not viable.
1.2 Motivation and contributions
[…]
2 Method
This section presents the missing value problem, incremental learning, the attention mask mechanism, and the proposed algorithm combining the attention mask within an incremental learning framework.
2.1 Preliminaries
A tabular data set $X\in\Re^{n\times d}=\{X_{obs},X_{miss}\}$ with $r\%$ missing values is composed of observed ($X_{obs}$) and missing ($X_{miss}$) values. An imputation model first initializes $X_{miss}$ using mean or median values as ($\hat{X}_{miss}$). The model then iteratively refines the initial estimates of missing values as $\hat{X}_{miss}^{i+1}$ using $X_{obs}$ and $\hat{X}_{miss}^{i}$ estimates from the previous iteration. An imputation model makes the assumption that $X_{obs}$ is predictive of $X_{miss}$, which may not always be true depending on the predictive model, data statistics, and missing value types. This assumption can be entirely waived when a data set can be learned using $X_{obs}$ alone without estimating the $X_{miss}$ portion. A data set with missing values can be learned through FIL using $X_{obs}$ alone. FIL updates an ML model when additional features become available in a multiphase data collection scenario. An ML model trained on the data set $X_{1}\in\Re^{N_{1}\times d_{1}}$ can be updated by a data set collected in the second phase $X_{2}\in\Re^{N_{2}\times(d_{1}+d_{2})}$ with additional $d_{2}$ features, and so on. Without FIL, the conventional approach combines the two tabular data sets $X_{1}$ and $X_{2}$ to train the ML model from scratch, where samples of $X_{1}$ would have values missing for additional $d_{2}$ features. Unfortunately, traditional ML is not usually designed to learn directly from data with missing values. In this context, we repurpose FIL to learn data with missing values directly, eliminating the need for data imputation or exclusion, as discussed in subsequent sections.
2.2 Proposed NIIL algorithm
The proposed No Imputation Incremental Learning (NIIL) completely avoids imputation or initialization of missing values $X_{miss}$. Suppose a labeled tabular data set with missing values $D=\{X,y\}$ has $n$ samples and $d$ features. First, we sort the features $f_{1},f_{2},\dots,f_{m}$ in ascending order of missing rates. The sorted features are partitioned into subsets of $k$ features with $s=ceil(\frac{k}{2})$ overlaps, as shown in Figure 1. The overlap between two consecutive feature subsets facilitates incremental learning using $P$ partitions of features, as follows.
$P=1+\lceil{\frac{d-k}{k-\lceil{k/2}\rceil}}\rceil$ (1)
These $P$ partitions are used to achieve incremental learning using a feature-tokenized transformer (FTT) [11, 33, 34] as follows. An FTT classifier model ($FTT_{1}$) is trained using the first partition $P_{1}=X\{f_{1},...,f_{k}\}$ of $k$ features. The next incremental learning session uses the second partition $P_{2}=X\{f_{s},...,f_{k+s}\}$ to train the previous $FTT_{1}$ model to $FTT_{2}$. Incremental learning continues until the $FTT_{P}$ model is obtained using the last feature partition, as illustrated in Figure 1. The trained $FTT_{P}$ model directly classifies a separate test data fold with missing values. Notably, FTT tokenizes features and uses a pre-trained language model to obtain embeddings for feature names and categorical features.In contrast, numerical values are passed through a linear projection layer to obtain corresponding embeddings. The tabular data representation is initialized by aggregating the embeddings of feature names and numerical and categorical feature values, which are subsequently fine-tuned for classification using a gated transformer. The proposed NIIL algorithm is presented in Algorithm 1. Missing values ($X^{miss}$) of X are excluded from attention scoring using attention masks designed as follows.
Algorithm 1 The NIIL Algorithm
  Input: Tabular data with missing values, $\{X\in\Re^{N\times d},y\}$
  Parameters: $k$: number of features in partitions.
  Output: Trained classifier model, $\mathcal{M}$ ($\theta$)
  Procedure:
  $\{X^{obs},X^{miss}\}\leftarrow\{X,y\}$
  Missing rate of $f_{j}$ feature, $r_{j}\leftarrow\frac{length(X^{miss}(:,j))}{length(X(:,j))}$
  $\{f_{1},f_{2},...,f_{d}\}\leftarrow$$\{r_{1}<r_{2}<...<r_{d}\}$, sorting $f_{i}$.
  $\{P_{1},P_{2},...\}$leftarrow $\{X_{1}(f_{1},f_{2}),X_{2}(f_{2},f_{3}),...\}$, when k=2.
  for  $P_{i}\in\{P_{1},P_{2},...\}$ do
     {$X_{i}^{obs},X_{i}^{miss}$} $\leftarrow P_{i}$
     $m\leftarrow X_{i}^{\text{miss}}$, Missing mask vector
     $M_{1}\leftarrow m$,  Eq. 2
     $M_{2}\leftarrow\bar{m}^{T}\bar{m}$
     $\mathcal{M}(\theta)\leftarrow\{(X_{i},M_{1},M_{2}),y_{i}\}$,  Eq. 5
  end for
A binary vector $m\in\Re^{1\times d}$ uses 1s to identify features with missing values and 0s otherwise. An outer product between a column vector of 1s and $m$ yields a $(d,d)$ binary matrix ($M_{1}$) with columns of 1s corresponding to the features with missing values. These columns of 1s are replaced with large negative values in Equation 2. An exponential function transforms large negative values in $M_{1}$ representing missing features to zeros, as shown in Equation 3.
$M_{1}$ $=$ $-\infty\odot(\mathbf{1}_{d}^{T}m)$ (2)
$exp(M_{1})$ $=$ $\begin{cases}0,&\text{if}\ M_{1}(:)=-\infty\\ 1,&\text{if}\ M_{1}(:)=0\end{cases}$ (3)
$M_{1}$ filters the attention columns corresponding to the features with missing values using the softmax function in Equation 2.2. The attention rows corresponding to features with missing values are eliminated using another binary mask, $M_{2}$. Here, $M_{2}=\bar{m}^{T}\bar{m}$ obtained from the complement of $m$ ($\bar{m}$) includes 0s in rows and columns that correspond to missing features and 1s otherwise.
Attention $=$ $\text{softmax}(\frac{QK^{T}}{\sqrt{h}}+M_{1})\odot M_{2}$
$=$ $(\frac{\exp{(\frac{QK^{T}}{\sqrt{h}}+M_{1}})}{\sum\exp{(\frac{QK^{T}}{\sqrt{h}}+M_{1}})}\odot M_{2})$
$=$ $\frac{\exp{(\frac{QK^{T}}{\sqrt{h}})*\exp{(M_{1})}}}{\sum\exp{(\frac{QK^{T}}{\sqrt{h}})*\exp{(M_{1})}}}\odot M_{2}$ (5)
Combining Equations 3 and 5, a $d\times h$ attention head in Equation 6 yields a $h$ dimensional embedding for each of the $d$ features. The query ($Q$), key ($K$), and value ($V$) matrices obtain the attention head in Equation 6.
Head ={($\text{Head =}\begin{cases}(0*I_{d\times h})*V,&\text{if $M_{1}(:)$\text{ = -$\infty$}, $M_{2}(:)=0$}\\ \text{softmax}(\frac{QK^{T}}{\sqrt{h}})V,&\text{if $M_{1}(:)=0$, $M_{2}(:)=1$}\end{cases}$ (6)
$M_{1}$ uses $\infty$- ∞ to mask the attention columns with missing features. The remaining rows of attention with missing features are masked by 0s of $M_{2}$. Otherwise, $M_{1}$ and $M_{2}$ retain the attention scores of all observed feature pairs using 0s and 1s, respectively.
3 Experiments
Table 1: Summary of the benchmark tabular data sets with heterogeneous data types and structures.
Type OpenML Data set Samples Features Numer- Categor Classes
Id ical -ical
Complete Data sets 13 Breast-cancer 286 9 0 9 2
31 Credit-g 1000 20 7 13 2
37 Diabetes 768 8 8 0 2
181 Yeast 1484 9 9 0 10
334 Monks-problems-2 601 7 0 7 2
463 Backache 180 31 5 26 2
1063 Kc2 522 21 21 0 2
1067 kc1 2109 21 21 0 2
1071 mw1 403 37 37 0 2
1480 Ilpd 583 10 9 1 2
1498 sa-heart 462 9 8 1 2
23381 Dresses-sales 500 12 1 11 2
40691 wine-quality-red 1599 11 11 0 6
Naturally Missing 55 Hepatitis 155 19 6 13 2
6332 Cylinder-bands 540 39 18 21 2
We simulate two types of missing values: 1) missing completely at random (MCAR) and 2) missing not at random (MNAR) at a missing rate between 10% and 50% at 10% intervals. Two data sets (IDs 55 and 6332) have naturally occurring missing values and are analyzed separately without simulating additional missing value types and rates.
3.1 Tabular data sets
Tabular data sets with large sample sizes ($>$10,000) are often selected to facilitate deep learning [35, 36]. However, most tabular data sets in practice have relatively smaller sample sizes ($<$1000) [37]. For generalizability, we have selected 15 diverse data sets with a variety of numerical and categorical feature mixtures from the OpenML [38] repository. The sample size of these data sets ranges from 155 to 2109, while the feature dimension is between 7 and 39. A summary of the tabular data sets used in this article is in Table 1, where the first 13 data sets are available in complete form without missing values. The remaining two data sets (IDs 55, and 6332) have missing values that occur naturally at the rate of 48.8% and 48.7%, respectively.
3.2 Model implementation and evaluation
The FTT classifier, adopted from [11], consists of a transformer with two encoder layers. The encoder layers transform numerical and categorical features into 128-dimensional embeddings. Each Transformer layer includes eight attention heads and a feedforward network with a 2048-dimensional hidden layer with ReLU activation functions and a dropout rate of 0.3. We have used the Adam optimizer with a fixed learning rate of $1\times 10^{-5}$ and no weight decay. The model is trained for up to 300 epochs using a batch size of 128. Early stopping is used with a patience of 50 epochs, terminating training when the validation loss ceases to improve for 50 consecutive epochs.
A five-fold cross-validation obtains the average area under the receiver operating characteristic curves (AUC) and the F1 scores. The average rank order of each method is reported in 15 tabular data sets based on AUC scores [8]. The five missing rates from 10% to 50% in 15 data sets yield 75 experimental scenarios for each missing value type. As in [39], we use a win matrix to compare methods in pairs by showing the percentage of experimental scenarios in which one method outperforms the other. The resilience of a model to varying missing rates is evaluated by comparing the AUC scores against the reference AUC score. The reference AUC score of a data set is obtained in its complete form before simulating missing values. We primarily evaluate the proposed incremental learning method for three partition sizes, k = 2, 3, and 4. Our analysis in Figure 5 suggests that FIL using feature partitions is computationally more efficient than learning the complete feature space ($d$ features) when k is less than $\lceil d/2\rceil$. Therefore, the performance of our model is reported for partitions when $k=\frac{d}{2}$.
3.3 Baseline methods
Our baseline methods have two types: one type requires imputation before classification, and the other can directly classify data without imputation. The requirement of imputation before classification is fulfilled by median value imputation, MICE [12], GAN-based imputation method GAIN [14], and diffusion-based imputation method, Diffputer [15]. Unlike mean values, median values are robust to outliers and are recommended for low missing rates. The subsequent classification of imputed data sets is achieved using two state-of-the-art models for tabular data, Gradient boosting trees (GBTs) and FTT [11]. On the other hand, XGBoost [27] and LightGBM [40] can use the Missing Incorporated in Attributes (MIA) strategy [28], as mentioned in Section 1.1, to directly classify data with missing values. XGBoost and LightGBM treat missing values as distinct categories during decision tree splits. The FTT classifier can directly classify missing values using an attention mask, presented in Equation 2.2. The proposed NIIL method is compared against 11 baseline methods: 1) median - GBT, 2) median - FTT, 3) MICE - GBT, 4) MICE - FTT, 5) GAIN - GBT, 6) GAIN - FTT, 7) Diffputer - GBT, 8) Diffputer - FTT, 9) AM - FTT, 10) MIA - XGBoost, and 11) MIA - LightGBM. All baseline methods are used with default hyperparameter settings.
4 Results
All results are obtained using an Ubuntu 22.04 machine utilizing an Intel(R) Xeon(R) W-2265 CPU (24 logical cores) running at 3.70GHz, 64GB of RAM, and a Quadro RTX A4000 GPU with 16GB of video memory. The performance of the proposed NIIL method is compared with the baseline methods using 1) the average rank based on AUC scores, 2) the win matrix for pairwise method comparison, and 3) the robustness of the performance to increasing missing rates. B presents the AUC scores specific to the 13 data sets of type MCAR in Table 6 and of type MNAR in Table 7.
Table 2: Average performance rank of the proposed (NIIL) and baseline models across 13 diverse MCAR-type data sets with five missing value rates.
Method Imputation 10% 20% 30% 40% 50% Avg. Rank Overall
Median-GBT Yes 7.8(0.05) 7.4(0.05) 7.5(0.04) 8.3(0.06) 7.2(0.05) 7.7(0.05) 7
Median-FTT Yes 8.7(0.07) 8.7(0.07) 8.9(0.06) 9.0(0.06) 9.8(0.07) 9.0(0.07) 12
MICE-GBT Yes 6.5(0.05) 6.8(0.05) 8.3(0.05) 8.2(0.05) 7.8(0.06) 7.5(0.05) 6
MICE-FTT Yes 8.0(0.05) 7.6(0.05) 6.9(0.04) 6.8(0.04) 7.2(0.05) 7.3(0.05) 5
GAIN-GBT Yes 10.0(0.05) 9.2(0.05) 7.4(0.05) 7.1(0.06) 6.6(0.04) 8.0(0.05) 8
GAIN-FTT Yes 7.8(0.06) 6.6(0.06) 6.8(0.06) 7.1(0.07) 6.8(0.06) 7.0(0.06) 4
Diffputer-GBT Yes 9.1(0.06) 9.1(0.06) 10.2(0.06) 11.9(0.06) 9.7(0.06) 10.0(0.06) 14
Diffputer-FTT Yes 8.7(0.06) 9.7(0.06) 10.0(0.06) 9.3(0.06) 10.1(0.06) 9.5(0.06) 13
MIA-Xgboost No 9.1(0.05) 8.9(0.05) 8.4(0.05) 8.9(0.05) 8.2(0.05) 8.7(0.05) 11
MIA-LightGBM No 8.5(0.04) 9.2(0.04) 9.0(0.05) 9.5(0.05) 7.2(0.05) 8.7(0.05) 10
AM-FTT No 9.3(0.06) 10.5(0.06) 10.8(0.06) 10.3(0.07) 11.5(0.07) 10.5(0.06) 15
NIIL (K =2) No 8.2(0.04) 8.2(0.05) 7.6(0.04) 7.8(0.05) 8.5(0.06) 8.0(0.05) 9
NIIL (K =3) No 6.7(0.05) 6.0(0.04) 5.9(0.04) 5.9(0.04) 6.2(0.05) 6.1(0.05) 2
NIIL (K =4) No 6.4(0.04) 6.0(0.04) 6.5(0.05) 5.2(0.04) 7.0(0.05) 6.2(0.05) 3
NIIL (K =d/2) No 5.5(0.05) 6.1(0.05) 5.8(0.05) 4.8(0.04) 6.2(0.05) 5.7(0.05) 1
Table 3: Average performance rank of the proposed (NIIL) and baseline models across 13 diverse MNAR-type data sets with five missing value rates.
Method Imputation 10% 20% 30% 40% 50% Avg. Rank Overall
Median-GBT Yes 7.2(0.05) 7.3(0.06) 7.3(0.06) 8.5(0.05) 5.8(0.06) 7.2(0.06) 6
Median-FTT Yes 7.4(0.05) 7.2(0.05) 6.5(0.06) 7.7(0.06) 8.8(0.06) 7.5(0.06) 9
MICE-GBT Yes 9.1(0.05) 8.0(0.04) 7.0(0.06) 7.2(0.06) 5.8(0.05) 7.4(0.05) 8
MICE-FTT Yes 11.1(0.05) 9.5(0.06) 10.1(0.06) 9.5(0.06) 10.0(0.07) 10.0(0.06) 13
GAIN-GBT Yes 8.5(0.04) 8.8(0.05) 9.9(0.04) 8.6(0.04) 8.8(0.04) 8.9(0.04) 12
GAIN-FTT Yes 11.5(0.07) 11.9(0.08) 11.2(0.06) 12.3(0.06) 12.9(0.06) 12.0(0.06) 15
Diffputer-GBT Yes 9.2(0.06) 10.5(0.07) 10.9(0.06) 10.0(0.05) 9.8(0.06) 10.1(0.06) 14
Diffputer-FTT Yes 8.4(0.07) 7.5(0.06) 8.8(0.07) 8.0(0.06) 10.1(0.06) 8.6(0.06) 11
MIA-Xgboost No 5.5(0.05) 6.8(0.05) 7.5(0.05) 8.5(0.05) 7.2(0.06) 7.1(0.05) 5
MIA-LightGBM No 6.5(0.05) 7.7(0.04) 7.2(0.05) 7.0(0.05) 7.9(0.06) 7.3(0.05) 7
AM-FTT No 9.9(0.06) 7.5(0.06) 8.2(0.05) 8.9(0.05) 8.2(0.06) 8.6(0.06) 10
NIIL (K =2) No 7.2(0.05) 7.1(0.06) 6.3(0.06) 6.4(0.05) 5.7(0.06) 6.5(0.06) 4
NIIL (K =3) No 6.6(0.05) 6.2(0.06) 6.8(0.07) 5.9(0.05) 6.2(0.06) 6.3(0.06) 2
NIIL (K =4) No 6.2(0.05) 7.2(0.06) 6.3(0.06) 5.7(0.05) 6.4(0.06) 6.4(0.06) 3
NIIL (K =d/2) No 5.8(0.04) 6.7(0.07) 6.1(0.06) 5.7(0.06) 6.4(0.06) 6.1(0.06) 1
Table 4: Mean AUC scores after classifying data sets (Hepatitis (ID: 55), Cylinder-bands (ID: 6332)) with naturally occurring missing values at the rate of 48.8% and 48.7%, respectively.
ID Median- Median- MICE- MICE- GAIN- GAIN- Diffputer- Diffputer- MIA- MIA- AM-FTT NIIL
GBT FTT GBT FTT GBT FTT GBT FTT Xgboost LightGBM FTT (K=d/2)
55 0.806 (0.08) 0.809 (0.04) 0.790 (0.09) 0.840 (0.04) 0.779 (0.12) 0.837 (0.05) 0.788 (0.10) 0.805 (0.06) 0.757 (0.06 0.782 (0.07) 0.806 (0.04) 0.846 (0.04)
6332 0.867 (0.02) 0.743 (0.01) 0.868 (0.02) 0.796 (0.04) 0.739 (0.07) 0.808 (0.03) 0.838 (0.03) 0.781 (0.05) 0.812 (0.01) 0.757 (0.06) 0.814 (0.03) 0.861 (0.03)
Rank 4.0 (0.05) 7.5 (0.02) 5.0 (0.05) 5.0 (0.04) 11.5 (0.10) 5.0 (0.04) 7.0 (0.07) 8.0 (0.06) 6.5 (0.03) 10.0 (0.07) 6.0 (0.04) 2.5 (0.03)
4.1 Average performance rank
Table 2 compares the proposed NIIL method with 11 baselines for MCAR-type data. In single-value imputation, median imputation is better when GBT is a classifier than FTT, regardless of missing rates. The MICE or multiple-imputation-based method outperforms median value imputation. The GAN-based imputation method (GAIN) with FTT performs better than MICE for all missing rates. In contrast, the diffusion-based imputation method performs the worst among all imputation techniques. Among the methods without explicit imputations, the MIA-based models (MIA-XGBoost and MIA-LightGBM) perform better than AM-FTT. The average performance rank shows that the proposed NIIL (K = d/2) outperforms all 11 baselines. NIIL with other feature partition sizes, including NIIL (K = 3) and NIIL (K = 4), also rank among the top three. Even at low missing rates (10%–30%), NIIL (K = d/2) maintains a strong performance rank, while at higher rates it demonstrates clear superiority. These results confirm the robustness and consistent performance of the proposed NIIL method for MCAR-type data.
Table 3 presents the average classification performance ranks in classifying MNAR-type data. Imputation-based methods that present competitive performance for MCAR-type data (GAIN-FTT and MICE-FTT) show a substantial performance drop for MNAR-type data, ranking 15th and 13th, respectively. In contrast, MIA-based models (MIA-XGBoost and MIA-LightGBM) perform better on MNAR data than MCAR data. This observation is consistent with previous findings [41] suggesting the MIA method performs better when the missingness is informative, such as MNAR. At lower missing rates (10% and 20%), the average ranks suggest that MIA methods outperform all methods that require an imputation step. However, median-GBT consistently outperforms MIA and other imputation-based methods when the missing rate is above 20%. On the other hand, NIIL with other partition sizes (K = 2, 3 and 4) is among the top four best methods, while (K = d / 2) is the best performing method. These results suggest that the proposed NIIL method is not impacted by missing value types, unlike other imputation-based approaches. Table 4 presents the performance of two data sets with naturally occurring missing values, Hepatitis (ID: 55) and Cylinder-bands (ID: 6332) with missing rates of 48.8% and 48.7%, respectively. The proposed NIIL method (K = d/2) achieves the best average performance rank (2.5 (0.03)), while the second best method is median-GBT (average rank 4.0 (0.05). This again shows the robustness of the NIIL method to missing value types.
(a) MCAR type
(b) MNAR type
Figure 2: Win matrix. Values are the fraction of experimental scenarios in which the row methods outperform methods in the columns.
4.2 Pairwise model comparison
We count the experimental scenarios in which one method wins over another. There are 65 experimental scenarios from 13 data sets and five missing rates.
Figure 2 (a) for the MCAR-type data shows that the proposed NIIL method is always superior to the baselines for more than 50% of the experimental scenarios. In the best case, NIIL outperforms Diffputer-FTT and MIA-Xgboost in 72% of the scenarios. Figure 2 (b) for the MNAR-type data shows that NIIL is always superior to the baselines for more than 50% of the experimental scenarios. In the best case, NIIL outperforms GAIN-FTT in 83% of the scenarios, which is 52% for MCAR-type data. This indicates that deep-imputation methods are not optimized for MNAR-type data.
4.3 Computational time and robustness to missing rates
(a) MCAR type
(b) MNAR type
Figure 3: Effect of increasing missing value rates. The average percentage of the reference AUC score is obtained across 13 data sets.
Table 5 presents the run time for the imputation and classification steps of the baseline methods. As expected, methods faster than NIIL include GAIN-GBT, Diffputer-GBT, AM-FTT, MIA-Xgboost, and MIA-LightGBM, mainly due to the fast GBT classifier model. However, the classification performance rank of these methods is worse than that of NIIL. One computational benefit of NIIL is that it saves the time required for imputation. Overall, the run-time is decent for tabular data sets of standard sample sizes (2̃000).
Figure 3 shows the effect of varying the missing rates on the performance of five top-performing models. The reference AUC scores for the robustness analysis are obtained from complete data sets with 0% missing values. The AUC score drops against the reference AUC due to increased missing rates. Figure 3(a) for MCAR-type data shows the robustness of the proposed method when the missing rate is higher than 30%. GAIN-FTT is the most robust method up to 30% missing values. However, Figure 3(b) for MNAR-type data shows that the robustness of the proposed NIIL method to high missing rates is better than all other baselines.
Table 5: Computational runtime in seconds for imputation (training and imputing) and classification (training and inference) steps using the credit-g data set with 50% MCAR type missing values.
Method Imputation Classification Total Classification Classification
runtime runtime runtime Rank (MCAR) Rank (MNAR)
MICE - GBT 417 2 419 6 7
MICE - FTT 417 190 607 5 13
GAIN - GBT 51 2 54 8 11
GAIN - FTT 51 190 241 3 15
Diffputer - GBT 8 1 9 13 14
Diffputer - FTT 8 190 198 12 10
AM-FTT N/A 21 21 15 12
MIA - Xgboost N/A 1 1 10 6
MIA - LightGBM N/A 1 1 9 8
NIIL (k = $\frac{d}{2}$) N/A 163 163 1 1
4.4 Ablation studies
In an ablation study, we vary the value of feature partition size k for incremental learning using three representative data sets: Kc2 (ID 1063), Diabetes (ID 37), and Dresses sales (ID 23381). Figure 4 shows the k values that yield the best AUC scores for individual data sets. In general, feature incremental learning is superior to learning with all features at a time (i.e., no incremental learning, k = d). The best average AUCs are obtained for the data sets Kc2 $(d=21)$ at k=10, Diabetes $(d=8)$ at k=6, and Dresses sales $(d=12)$ at k = 6. Therefore, the value of k can be chosen as half of the feature dimension for the best classification performance. This observation aligns with the computational cost analysis presented in Figure 5 in  A, where $k<\frac{d}{2}$ is computationally more efficient than using the entire feature space ($d$).NIIL outperforms all or most of the baselines at k = 3, which can be further improved by increasing the partition size to $\frac{d}{2}$. The proposed method is evaluated on tabular data sets with categorical-only, numerical-only, and mixed data types. The sample size of these data sets ranges from 155 to 2109. Most tabular data in practice have limited sample sizes ($<1000$) with heterogeneous data types. The imputation of data with limited samples poses an additional challenge. In contrast, deep learning studies selectively use tabular data sets with large sample sizes ($>10000$) only [35, 8]. Therefore, our results are presented using the most practical tabular data sets, missing types, and rates to ensure generalizability.
5 Discussions
The paper proposes incremental learning for tabular data to exclude the requirement of imputing missing values. The findings of the paper can be summarized as follows. First, the proposed NIIL method outperforms all baseline methods with and without imputation regardless of the missing value types. The exclusion of missing value imputation steps makes the proposed model insensitive to missing value types, unlike other imputation-based approaches. Second, deep imputation methods may not be optimally designed to impute all missing value types, including the MNAR-type. The proposed NIIL method is a superior choice when the missing value type is unknown or complex. Third, the attention mask enables the classification of missing values without imputation. However, the performance of the AM-FTT model shows that the attention mask itself or the FTT model does not improve the classification performance. Incremental learning plays an important role along with AM-FTT to enable the proposed NIIL method to outperform the baselines. Fourth, the NIIL method offers the best performance and computational time trade-off with better resilience to missing value rates than baselines.
Figure 4: Effects of partition size (k) on average AUC scores obtained across varying missing value rates. The total number of features ($d$) for three data sets are: the Kc2 data set ($d=21$), Diabetes data set ($d=8$), and Dresses Sales data set ($d=12$).
5.1 Robustness to missing value types
The proposed method brings several advantages over existing imputation methods. First, imputation methods are usually not designed to consider and customize for different types of missing values. Therefore, an effective imputation method for MCAR-type data may not be optimal for MNAR-type data or data that naturally come with missing values. The imputation error and variance in the data due to varying missing value types can be avoided using our non-imputation classification method. Moreover, NIIL maximizes the use of observed values in learning, preserving data integrity. This characteristic is particularly valuable in sensitive domains such as healthcare and scientific research, where model-generated synthetic values are often considered untrustworthy and can compromise the credibility of data-driven outcomes. Second, imputation of heterogeneous feature space in tabular data (e.g. numerical, categorical) requires different learning frameworks. For example, continuous numerical value imputation is a regression problem, whereas categorical variables need classification models. The proposed method alleviates the need for separate computational steps and the requirement of train-test splits for the imputation process.
5.2 Robustness to high missing rates
Imputation-based methods typically learn missing values from their relationship with observed values. However, as the missing value rate increases, the relationship ceases to exist, causing a decline in imputation quality and downstream predictive accuracy. Many existing studies evaluate imputation methods up to a 30% missing rate [42, 43]. The robustness of many similar methods to higher missing rates remains unknown. Real-world scenarios may include important features missing in more than 50% of the samples, as observed in [44]. Imputation methods suffer from higher error rates, resulting in more unreliable data for downstream classification. However, methods without requiring imputation, such as AM-FTT, rely solely on observed values. The proposed NIIL approach avoids the imputation step and instead leverages feature attention via the FTT framework and incremental learning of observed values to remain robust against increasing missing value rates.
5.3 Comparison with the literature
The literature explores a variety of DL/ML methods for missing-value imputation. A recent survey found that deep generative models (e.g., GAIN, VAE) require a large sample size ($n>30,000$); otherwise, traditional imputation methods (e.g., MICE) tend to outperform generative approaches [18]. Similar requirements for the performance of imputation methods do not apply to NIIL. This may explain why MICE is superior to GAIN in our results for tabular data sets with a standard sample size. Furthermore, the recent diffusion-based method, Diffputer [15] provided a comprehensive benchmarking against multiple advanced baselines, including diffusion-based methods (MissDiff [21], TabCSDI [22]), generative adversarial methods (GAIN [14], MIWAE [45]), iterative imputation (MICE[46], EM, HyperImpute [19]), and graph neural network-based approaches (MIRACLE [47], GRAPE [48]). Diffputer demonstrates better performance than these baseline methods, particularly highlighting its suitability for large datasets ($n>10,000$). Our results suggest that Diffputer performance may not generalize well across different data sets, missing value rates, and types, unlike the proposed NIIL method. It may be inferred that NIIL would also outperform those advanced baseline deep learning methods for imputing missing values. Despite the advantages of deep imputation methods in large-scale datasets, practitioners must carefully consider the missing value rates, types, and data characteristics when selecting appropriate imputation methods for real-world applications.
5.4 Broader impact
This paper introduces a novel machine learning method for tabular data with missing values. Missing values appear in varying rates and types in many application domains, where one imputation method may not be optimal for all scenarios. Missing values in tabular data sets critically challenge the feasibility and performance of machine learning models. The widespread adoption of arbitrary missing value imputation strategies raises many questions about data quality and information reliability among data stakeholders. Our proposed ML method enables machine learning without missing value imputation, alleviating data wastage, biases, errors, and computational costs associated with missing value imputations. The proposed method foresees positive social impacts through increased inclusivity, integrity, trustworthiness, and efficiency in data analysis workflows. The impact on health science and the analysis of health records data is significant, where missing values are a major barrier to the development of data-driven precision medicine. However, practitioners should consider ethical implications, such as the risks of reinforcing biases in the original data set or overlooking the reasons behind missing data.
5.5 Limitations
The proposed NIIL method performs remarkably in deep learning tabular data without imputation or initializing missing values. Although deep learning without imputation of missing values is a significant contribution, the proposed method has several limitations. When the missing value rate is low and completely random, sophisticated methods for imputing missing values can yield better results than our proposed non-imputation method. Several imputation and classification models are computationally more efficient than the proposed non-imputation classification method. However, the difference in runtime is around several minutes for tabular data sets of standard sizes. When data quality and trustworthiness matter, especially in the most practical missing value scenarios (naturally occurring and MNAR type), the proposed NIIL method results in the overall best performance.
6 Conclusions
This paper introduces NIIL, a novel incremental learning algorithm that is performance-wise effective, robust to missing value rates and types, and computationally efficient for learning tabular data with missing values. Although attention masks in the feature-tokenized transformer enable no imputation learning of tabular data with missing values, incremental learning additionally yields substantial and consistent performance improvement. The proposed NIIL algorithm outperforms state-of-the-art deep imputation methods in downstream classification tasks, in addition to its ability to learn the high-dimensional feature space incrementally. Empirically, we show that the most computationally efficient and performance-wise effective feature partition size for incremental learning is half of the feature dimension of the data set.
7 Acknowledgments
Appendix A Number of operations and computational cost
Figure 5: Effects of feature partition size (k) on the computational cost of incremental learning compared to learning the complete feature space (d) at a time. The vertical dotted lines mark the feature partition size when its computational cost equals that of learning the complete feature space or the relative operation ratio is 1.0).
Appendix B Data set-specific AUC classification scores
Table 6: AUC value for MCAR missing type for two missing rates: 10% and 50%. Rows are data sets, columns are the models.
Dataset ID % Median-GBT Median-FTT MICE-GBT MICE-FTT GAIN-GBT GAIN-FTT Diffputer-FTT Diffputer-GBT MIA-Xgboost MIA-LightGBM AM-FTT NIIL (K=d/2)
1063 10 0.840 (0.06) 0.840 (0.05) 0.837 (0.04) 0.838 (0.04) 0.795 (0.04) 0.846 (0.04) 0.847 (0.04) 0.820 (0.05) 0.791 (0.05) 0.795 (0.07) 0.856 (0.04) 0.862 (0.04)
50 0.802 (0.05) 0.732 (0.02) 0.824 (0.06) 0.833 (0.04) 0.772 (0.06) 0.836 (0.05) 0.849 (0.05) 0.785 (0.08) 0.780 (0.05) 0.773 (0.05) 0.860 (0.04) 0.854 (0.04)
463 10 0.642 (0.16) 0.717 (0.21) 0.653 (0.06) 0.694 (0.09) 0.672 (0.11) 0.617 (0.24) 0.566 (0.21) 0.654 (0.12) 0.634 (0.08) 0.630 (0.06) 0.587 (0.25) 0.728 (0.05)
50 0.664 (0.09) 0.621 (0.21) 0.565 (0.09) 0.714 (0.11) 0.601 (0.03) 0.596 (0.27) 0.481 (0.18) 0.593 (0.13) 0.583 (0.13) 0.501 (0.11) 0.523 (0.14) 0.674 (0.10)
13 10 0.713 (0.07) 0.505 (0.10) 0.713 (0.07) 0.691 (0.05) 0.659 (0.07) 0.704 (0.06) 0.647 (0.07) 0.670 (0.07) 0.628 (0.08) 0.643 (0.04) 0.661 (0.07) 0.660 (0.07)
50 0.672 (0.09) 0.489 (0.08) 0.672 (0.09) 0.623 (0.04) 0.628 (0.05) 0.652 (0.07) 0.561 (0.08) 0.628 (0.05) 0.588 (0.10) 0.629 (0.06) 0.488 (0.10) 0.571 (0.11)
31 10 0.761 (0.03) 0.747 (0.03) 0.749 (0.04) 0.756 (0.03) 0.737 (0.03) 0.733 (0.02) 0.761 (0.04) 0.743 (0.02) 0.743 (0.03) 0.680 (0.04) 0.746 (0.03) 0.753 (0.03)
50 0.657 (0.03) 0.658 (0.06) 0.646 (0.04) 0.704 (0.04) 0.705 (0.04) 0.693 (0.05) 0.695 (0.03) 0.657 (0.03) 0.666 (0.05) 0.642 (0.03) 0.656 (0.05) 0.674 (0.04)
37 10 0.836 (0.03) 0.841 (0.04) 0.841 (0.03) 0.772 (0.04) 0.794 (0.02) 0.803 (0.03) 0.804 (0.03) 0.805 (0.03) 0.812 (0.03) 0.823 (0.03) 0.804 (0.03) 0.804 (0.03)
50 0.754 (0.05) 0.763 (0.06) 0.724 (0.05) 0.687 (0.05) 0.737 (0.02) 0.548 (0.10) 0.683 (0.06) 0.696 (0.07) 0.706 (0.04) 0.716 (0.04) 0.703 (0.04) 0.718 (0.05)
23381 10 0.577 (0.05) 0.580 (0.03) 0.588 (0.05) 0.582 (0.03) 0.621 (0.04) 0.576 (0.09) 0.593 (0.06) 0.570 (0.05) 0.530 (0.07) 0.557 (0.05) 0.603 (0.02) 0.623 (0.03)
50 0.523 (0.06) 0.488 (0.10) 0.520 (0.06) 0.488 (0.10) 0.527 (0.05) 0.561 (0.04) 0.557 (0.05) 0.560 (0.09) 0.509 (0.03) 0.569 (0.04) 0.476 (0.06) 0.516 (0.03)
1480 10 0.727 (0.03) 0.708 (0.04) 0.734 (0.05) 0.742 (0.04) 0.731 (0.02) 0.724 (0.05) 0.723 (0.07) 0.733 (0.03) 0.720 (0.05) 0.738 (0.05) 0.663 (0.04) 0.729 (0.05)
50 0.715 (0.07) 0.620 (0.09) 0.684 (0.07) 0.696 (0.02) 0.679 (0.03) 0.650 (0.05) 0.641 (0.06) 0.680 (0.03) 0.665 (0.04) 0.669 (0.05) 0.596 (0.06) 0.678 (0.06)
1067 10 0.800 (0.02) 0.791 (0.02) 0.810 (0.02) 0.793 (0.02) 0.799 (0.02) 0.796 (0.02) 0.807 (0.02) 0.806 (0.03) 0.992 (0.01) 0.993 (0.00) 0.797 (0.02) 0.875 (0.16)
50 0.778 (0.02) 0.695 (0.03) 0.792 (0.03) 0.781 (0.02) 0.784 (0.03) 0.788 (0.01) 0.785 (0.02) 0.771 (0.03) 0.796 (0.04) 0.801 (0.03) 0.724 (0.02) 0.793 (0.02)
334 10 0.684 (0.07) 0.721 (0.13) 0.684 (0.07) 0.805 (0.12) 0.628 (0.09) 0.856 (0.01) 0.721 (0.13) 0.717 (0.08) 0.803 (0.04) 0.756 (0.06) 0.822 (0.11) 0.853 (0.08)
50 0.528 (0.06) 0.531 (0.04) 0.528 (0.06) 0.565 (0.04) 0.537 (0.09) 0.577 (0.07) 0.531 (0.04) 0.532 (0.04) 0.603 (0.01) 0.620 (0.04) 0.553 (0.02) 0.598 (0.07)
1071 10 0.650 (0.10) 0.682 (0.19) 0.778 (0.07) 0.778 (0.08) 0.657 (0.11) 0.795 (0.06) 0.734 (0.10) 0.710 (0.12) 0.762 (0.05) 0.760 (0.03) 0.712 (0.15) 0.798 (0.03)
50 0.695 (0.06) 0.665 (0.21) 0.690 (0.08) 0.746 (0.11) 0.685 (0.07) 0.767 (0.09) 0.741 (0.09) 0.650 (0.07) 0.650 (0.09) 0.676 (0.08) 0.667 (0.15) 0.742 (0.06)
1498 10 0.700 (0.03) 0.755 (0.03) 0.701 (0.03) 0.744 (0.04) 0.698 (0.06) 0.745 (0.05) 0.738 (0.03) 0.704 (0.05) 0.671 (0.04) 0.675 (0.03) 0.733 (0.04) 0.738 (0.04)
50 0.656 (0.03) 0.693 (0.03) 0.668 (0.05) 0.731 (0.04) 0.683 (0.03) 0.707 (0.03) 0.649 (0.06) 0.684 (0.06) 0.631 (0.06) 0.641 (0.05) 0.654 (0.09) 0.652 (0.08)
40691 10 0.770 (0.03) 0.746 (0.01) 0.789 (0.03) 0.704 (0.04) 0.758 (0.05) 0.735 (0.02) 0.733 (0.02) 0.736 (0.05) 0.834 (0.02) 0.829 (0.02) 0.724 (0.02) 0.744 (0.02)
50 0.692 (0.04) 0.669 (0.01) 0.685 (0.04) 0.644 (0.02) 0.698 (0.01) 0.682 (0.02) 0.604 (0.02) 0.604 (0.05) 0.699 (0.03) 0.718 (0.04) 0.654 (0.02) 0.664 (0.01)
181 10 0.819 (0.04) 0.792 (0.02) 0.798 (0.05) 0.744 (0.02) 0.803 (0.03) 0.808 (0.02) 0.784 (0.02) 0.803 (0.04) 0.832 (0.03) 0.835 (0.03) 0.738 (0.01) 0.768 (0.01)
50 0.686 (0.06) 0.716 (0.02) 0.701 (0.06) 0.642 (0.02) 0.710 (0.05) 0.505 (0.02) 0.634 (0.02) 0.698 (0.06) 0.731 (0.05) 0.724 (0.05) 0.642 (0.02) 0.704 (0.01)
Table 7: AUC value for MNAR missing type for two missing rates: 10% and 50%. Rows are data sets, columns are the models.
Dataset ID % Median-GBT Median-FTT MICE-GBT MICE-FTT GAIN-GBT GAIN-FTT Diffputer-GBT Diffputer-FTT MIA-Xgboost MIA-LightGBM AM-FTT NIIL (K=d/2)
1063 10 0.892 (0.05) 0.623 (0.11) 0.863 (0.04) 0.668 (0.14) 0.496 (0.03) 0.509 (0.02) 0.468 (0.06) 0.525 (0.11) 0.904 (0.04) 0.912 (0.04) 0.783 (0.07) 0.871 (0.05)
50 0.553 (0.06) 0.536 (0.05) 0.506 (0.04) 0.499 (0.05) 0.505 (0.02) 0.512 (0.03) 0.507 (0.08) 0.486 (0.05) 0.540 (0.07) 0.542 (0.08) 0.492 (0.04) 0.557 (0.03)
463 10 0.627 (0.09) 0.675 (0.06) 0.596 (0.11) 0.485 (0.12) 0.672 (0.10) 0.627 (0.22) 0.532 (0.08) 0.541 (0.21) 0.665 (0.08) 0.595 (0.10) 0.406 (0.09) 0.746 (0.06)
50 0.652 (0.13) 0.595 (0.12) 0.618 (0.14) 0.439 (0.12) 0.557 (0.09) 0.595 (0.15) 0.663 (0.09) 0.563 (0.11) 0.623 (0.10) 0.543 (0.11) 0.602 (0.21) 0.657 (0.10)
13 10 0.704 (0.08) 0.657 (0.08) 0.704 (0.08) 0.645 (0.06) 0.701 (0.09) 0.527 (0.13) 0.666 (0.08) 0.694 (0.06) 0.698 (0.08) 0.655 (0.06) 0.657 (0.08) 0.676 (0.07)
50 0.642 (0.09) 0.583 (0.08) 0.642 (0.09) 0.630 (0.11) 0.676 (0.10) 0.487 (0.16) 0.621 (0.07) 0.674 (0.09) 0.602 (0.10) 0.638 (0.10) 0.583 (0.08) 0.607 (0.10)
31 10 0.696 (0.01) 0.745 (0.04) 0.690 (0.01) 0.534 (0.06) 0.730 (0.03) 0.537 (0.11) 0.710 (0.04) 0.739 (0.04) 0.693 (0.03) 0.680 (0.04) 0.494 (0.02) 0.742 (0.03)
50 0.637 (0.02) 0.679 (0.01) 0.629 (0.03) 0.540 (0.06) 0.657 (0.03) 0.507 (0.02) 0.604 (0.06) 0.643 (0.04) 0.661 (0.03) 0.642 (0.03) 0.531 (0.03) 0.684 (0.02)
37 10 0.510 (0.02) 0.525 (0.03) 0.491 (0.05) 0.524 (0.04) 0.497 (0.01) 0.502 (0.01) 0.554 (0.05) 0.522 (0.02) 0.540 (0.02) 0.530 (0.02) 0.498 (0.03) 0.503 (0.04)
50 0.504 (0.05) 0.489 (0.07) 0.527 (0.05) 0.510 (0.04) 0.502 (0.02) 0.500 (0.01) 0.537 (0.05) 0.488 (0.05) 0.502 (0.04) 0.526 (0.03) 0.494 (0.04) 0.502 (0.04)
23381 10 0.570 (0.06) 0.625 (0.06) 0.570 (0.06) 0.486 (0.02) 0.557 (0.04) 0.538 (0.08) 0.531 (0.16) 0.621 (0.06) 0.555 (0.05) 0.582 (0.05) 0.557 (0.10) 0.615 (0.04)
50 0.581 (0.06) 0.573 (0.05) 0.586 (0.07) 0.506 (0.01) 0.566 (0.13) 0.452 (0.11) 0.435 (0.15) 0.543 (0.02) 0.599 (0.03) 0.499 (0.04) 0.551 (0.06) 0.566 (0.05)
1480 10 0.542 (0.05) 0.534 (0.04) 0.539 (0.04) 0.520 (0.04) 0.547 (0.02) 0.511 (0.05) 0.581 (0.05) 0.534 (0.03) 0.558 (0.03) 0.551 (0.04) 0.518 (0.07) 0.559 (0.06)
50 0.537 (0.07) 0.543 (0.04) 0.531 (0.03) 0.528 (0.04) 0.556 (0.03) 0.479 (0.04) 0.546 (0.05) 0.530 (0.06) 0.519 (0.08) 0.523 (0.08) 0.535 (0.04) 0.546 (0.01)
1067 10 0.802 (0.02) 0.690 (0.04) 0.987 (0.01) 0.957 (0.01) 0.503 (0.01) 0.462 (0.09) 0.486 (0.01) 0.798 (0.11) 0.992 (0.01) 0.993 (0.01) 0.978 (0.01) 0.880 (0.04)
50 0.701 (0.08) 0.600 (0.02) 0.796 (0.03) 0.732 (0.06) 0.510 (0.01) 0.550 (0.09) 0.484 (0.03) 0.568 (0.04) 0.796 (0.04) 0.801 (0.03) 0.792 (0.02) 0.647 (0.09)
334 10 0.643 (0.05) 0.806 (0.10) 0.643 (0.05) 0.767 (0.11) 0.588 (0.05) 0.499 (0.04) 0.622 (0.06) 0.742 (0.17) 0.811 (0.04) 0.767 (0.04) 0.806 (0.10) 0.843 (0.02)
50 0.510 (0.01) 0.575 (0.05) 0.510 (0.01) 0.537 (0.05) 0.504 (0.06) 0.481 (0.03) 0.495 (0.04) 0.510 (0.04) 0.578 (0.02) 0.553 (0.05) 0.575 (0.05) 0.621 (0.07)
1071 10 0.584 (0.08) 0.462 (0.10) 0.526 (0.06) 0.464 (0.10) 0.525 (0.03) 0.483 (0.03) 0.470 (0.04) 0.526 (0.09) 0.536 (0.07) 0.546 (0.07) 0.466 (0.12) 0.430 (0.05)
50 0.612 (0.03) 0.421 (0.11) 0.547 (0.12) 0.442 (0.12) 0.508 (0.01) 0.503 (0.01) 0.448 (0.08) 0.404 (0.06) 0.468 (0.11) 0.445 (0.10) 0.463 (0.07) 0.429 (0.16)
1498 10 0.616 (0.09) 0.633 (0.05) 0.593 (0.09) 0.484 (0.03) 0.632 (0.06) 0.614 (0.10) 0.571 (0.05) 0.617 (0.06) 0.546 (0.05) 0.572 (0.06) 0.512 (0.04) 0.642 (0.06)
50 0.569 (0.06) 0.598 (0.05) 0.556 (0.03) 0.431 (0.06) 0.577 (0.02) 0.457 (0.06) 0.576 (0.05) 0.587 (0.03) 0.536 (0.09) 0.530 (0.08) 0.543 (0.07) 0.602 (0.04)
40691 10 0.555 (0.04) 0.486 (0.03) 0.488 (0.06) 0.497 (0.02) 0.502 (0.03) 0.497 (0.01) 0.500 (0.02) 0.499 (0.03) 0.537 (0.04) 0.539 (0.05) 0.713 (0.01) 0.478 (0.04)
50 0.512 (0.05) 0.506 (0.02) 0.511 (0.01) 0.505 (0.03) 0.500 (0.02) 0.487 (0.01) 0.490 (0.03) 0.499 (0.03) 0.474 (0.06) 0.490 (0.05) 0.642 (0.02) 0.494 (0.03)
181 10 0.524 (0.03) 0.514 (0.01) 0.531 (0.03) 0.521 (0.02) 0.502 (0.04) 0.505 (0.02) 0.510 (0.05) 0.514 (0.01) 0.564 (0.03) 0.571 (0.04) 0.575 (0.01) 0.563 (0.02)
50 0.473 (0.02) 0.507 (0.02) 0.493 (0.02) 0.507 (0.02) 0.502 (0.05) 0.499 (0.01) 0.505 (0.03) 0.496 (0.02) 0.528 (0.03) 0.521 (0.02) 0.522 (0.01) 0.511 (0.02)